# OSDFace: One-Step Diffusion Model for Face Restoration

### v1

@@ 1 Introduction
Figure 1: Visual samples of diffusion-based face restoration methods. We provide multiply-accumulate operations (MACs), time, and number of timesteps during inference. OSEDiff* [58] is retrained on FFHQ dataset. Our OSDFace achieves a more natural and faithful visual result than other ones.
@@
As shown in Fig. 1, most diffusion-based face restoration methods can effectively recover basic facial features such as eyes, mouth, and contour. However, details, such as realistic-looking hair and complex background, often remain unharmonious. The primary issue lies in the insufficient incorporation of face priors. Some diffusion-based methods [68, 31, 6] rely solely on the diffusion model without considering face priors, resulting in unrealistic restorations. Other methods [62, 48] attempt to use face priors, but either limit the generative capabilities or cause information reduction. On the other hand, existing general OSD image restoration models, such as OSEDiff [58], are not specifically designed for facial features. OSEDiff employs image-to-tag prior and works well for natural images. However, it meets challenges when applied to face images, as humans are highly sensitive to facial features, and even minor inconsistencies or unrealistic details are easily noticeable.
@@ 2.2 Diffusion Models
Reducing the inference steps of diffusion models is critical to accelerating generation speeds and reducing computational costs. Yet, further reduction often results in a significant performance drop, making it crucial to balance inference speed with model capability. Most one-step diffusion (OSD) methods employ distillation to learn from a teacher model, ensuring the quality of the generated images [66, 65, 55, 58]. DFOSD [27] achieves promising results with OSD, free from the limitations of model distillation. However, all models are designed primarily for natural image restoration and lack strong generalization for face-specific tasks. It is essential to incorporate prior knowledge specific to human faces to enhance the applicability of OSD models in face restoration. Consequently, the models could generate more realistic and high-fidelity face images.
@@ 3.1 One-Step Diffusion (OSD) Model
Unlike one-step T2I diffusion models [47, 66], the UNet input in OSDFace is not entirely Gaussian noise. Therefore, we predefine a parameter $T_{L}\in[0,T]$ and feed it into the UNet, where $T$ represents the total diffusion timesteps (in Stable Diffusion (SD), $T$ is set to 1,000). Finally, we use the decoder $D_{\theta}$ to reconstruct the HQ face $\hat{I}_{H}$ from $\hat{z}_{H}$:
@@
In addition, the accurate generation of edge details is essential for achieving perceived sharpness in face restoration. Thus, following the practice in DFOSD [27], we introduce an edge-aware DISTS (EA-DISTS) as our perceptual loss
@@
One-Step Diffusion OSEDiff [58] (s=1) 0.4708 0.3165 51.2566 6.6968 72.7112 10.9948 89.8160 68.3579
@@
Table 1: Quantitative comparison on the synthetic CelebA-Test dataset with non-diffusion, multi-step diffusion, and one-step diffusion methods. The best and second best results are colored with red and blue. OSEDiff* is retrained on the FFHQ dataset for reference.
4 Experiments
@@ 4.1 Experimental Settings
Metrics. For the Synthetic Dataset, we employ LPIPS [70] and DISTS [9] as reference-based perceptual quality measures, together with MUSIQ [24] and NIQE [69] as no-reference image quality measures. To evaluate the distribution similarity between the ground truth and restored faces, we calculate FID [16] with both FFHQ and CelebA-Test HQ, showing the ability to recover “real” faces and HQ images. Following previous works [53, 15, 50], we use the embedding angle of ArcFace [8], namely ‘Deg.’, and landmark distance, namely ‘LMD’, as additional identity and fidelity metrics. For the Real-world Datasets, we employ non-reference metrics, including FID (with FFHQ as reference), as well as CLIPIQA [52], MANIQA [63], MUSIQ [24], and NIQE [69]. We utilize the evaluation codes provided by pyiqa [4] and VQFR [15] for all metrics.
@@
In the second stage, our OSDFace is trained by the AdamW optimizer [33], with learning rate 1×$\times 10$-4 and batch size 2 for both the generator and discriminator. The LoRA rank in the UNet is set to 16. The SD 2.1-base [1] serves as the pretrained OSD model, with the VRE used to construct the prompt embeddings. The setting of the discriminator is followed by DFOSD [27], and the generator and discriminator are trained alternately. Training is performed for 150K iterations on 2 NVIDIA A6000 GPUs.
Compared State-of-the-Art (SOTA) Methods. We compare OSDFace with several SOTA methods, including RestoreFormer++ [57], VQFR [15], CodeFormer [72], DAEFR [50], PGDiff [62], DifFace [68] and DiffBIR [31]. We retrain the current SOTA one-step diffusion model OSEDiff [58] on the FFHQ dataset, namely OSEDiff*.
@@ 4.2 Performance on Synthetic Data
Quantitative Results. The results in Tab. 1 show that OSDFace outperforms other methods in Deg., indicating a closer alignment with the identity features of HQ images. Our method also achieves obviously higher scores on LPIPS and DISTS, demonstrating superior perceptual coherence and detail preservation ability. The performance on FID (HQ) shows that OSDFace could excellently recover the distribution of original HQ dataset.
@@ 4.3 Performance on Real-world Data
Qualitative Results. Figure 6 visualizes representative images in three different real-world datasets. Existing face restoration methods often emphasize facial features while overlooking the precise reconstruction of facial accessories, such as glasses, hair, hats, headwear, and background details. Some methods, especially DiffBIR [31], are over-smoothing and lack fine facial textures. Our OSDFace can naturally and clearly restore multiple faces within a single image, demonstrating its effectiveness for unaligned faces. OSDFace ensures high-fidelity restoration with balanced and consistent visual quality throughout the entire image.

### v2

@@ 1 Introduction
Figure 1: Visual samples of diffusion-based face restoration methods. We provide multiply-accumulate operations (MACs), time, and number of timesteps during inference. Our OSDFace achieves a more natural and faithful visual result than other ones.
@@
As shown in Fig. 1, most diffusion-based face restoration methods can effectively recover basic facial features such as eyes, mouth, and contour. However, details, such as realistic-looking hair and complex backgrounds, often remain unharmonious. The primary issue lies in the insufficient incorporation of face priors [62]. Some diffusion-based methods [70, 32, 7] rely solely on the diffusion model without considering face priors, resulting in unrealistic restorations. Other methods [64, 49] attempt to use face priors, but either limit the generative capabilities or cause information reduction. On the other hand, existing general OSD image restoration models, such as OSEDiff [59], are not specifically designed for facial features. OSEDiff employs image-to-tag prior and works well for natural images. However, it meets challenges when applied to faces, as humans are highly sensitive to facial features, and even minor inconsistencies or unrealistic details are easily noticeable.
@@ 2.2 Diffusion Models
Reducing the inference steps of diffusion models is critical to accelerating generation speeds and reducing computational costs. Yet, further reduction often results in a significant performance drop, making it crucial to balance inference speed with model capability. Most one-step diffusion (OSD) methods employ distillation to learn from a teacher model, ensuring the quality of the generated images [68, 67, 56, 59]. D3SR [28] achieves promising results with OSD, free from the limitations of model distillation. However, all models are designed primarily for natural image restoration and lack strong generalization for face-specific tasks. It is essential to incorporate prior knowledge specific to human faces to enhance the applicability of OSD models in face restoration. Consequently, the models could generate more realistic and high-fidelity face images.
@@ 3.1 One-Step Diffusion (OSD) Model
Unlike one-step T2I diffusion models [48, 68], the UNet input in OSDFace is not entirely Gaussian noise. Therefore, we predefine a parameter $T_{L}\in[0,T]$ and feed it into the UNet, where $T$ represents the total diffusion timesteps (in Stable Diffusion, $T$ is set to 1,000). Finally, we use the decoder $D_{\theta}$ to reconstruct the HQ face $\hat{I}_{H}$ from $\hat{z}_{H}$:
@@
In addition, the accurate generation of edge details is essential for achieving perceived sharpness in face restoration. Thus, following the practice in D3SR [28], we introduce an edge-aware DISTS (EA-DISTS) as our perceptual loss
@@
One-Step Diffusion OSEDiff [59] (s=1) 0.3306 0.2170 71.4674 5.1241 67.3904 6.4141 73.4843 37.2103
@@
Table 1: Quantitative comparison on the synthetic CelebA-Test dataset with non-diffusion, multi-step diffusion, and one-step diffusion methods. The best and second best results are colored with red and blue.
Refer to caption Refer to caption LQ OSDFace Refer to caption Refer to caption LQ OSDFace
Figure 6: Zero-shot results of real-world cartoons.
@@ 4.1 Experimental Settings
Metrics. For the Synthetic Dataset, we employ LPIPS [72] and DISTS [10] as reference-based perceptual quality measures, together with MUSIQ [25] and NIQE [71] as no-reference image quality measures. To evaluate the distribution similarity between the ground truth and restored faces, we calculate FID [17] with both FFHQ and CelebA-Test HQ, showing the ability to recover “real” faces and HQ images. Following previous works [54, 16, 51], we use the embedding angle of ArcFace [9], namely “Deg.”, and landmark distance, namely “LMD”, as additional identity and fidelity metrics. For the Real-world Datasets, we employ non-reference metrics, including FID (with FFHQ as reference), as well as CLIPIQA [53], MANIQA [65], MUSIQ [25], and NIQE [71]. We utilize the evaluation codes provided by pyiqa [5] and VQFR [16] for all metrics.
@@
In the second stage, our OSDFace is trained by the AdamW optimizer [34], with learning rate $1$×$\times 10$-4 and batch size $2$ for both the generator and discriminator. The LoRA rank in the UNet is set to $16$. The SD 2.1-base [1] serves as the pretrained OSD model, with the VRE used to construct the prompt embeddings. The setting of the discriminator is followed by D3SR [28], and the generator and discriminator are trained alternately. Training is performed for 150K iterations on 2 NVIDIA A6000 GPUs.
Compared State-of-the-Art (SOTA) Methods. We compare OSDFace with several SOTA methods, including RestoreFormer++ [58], VQFR [16], CodeFormer [74], DAEFR [51], PGDiff [64], DifFace [70] and DiffBIR [32]. We also provide the current SOTA OSD model OSEDiff [59], and retrain it on FFHQ [23], namely OSEDiff*.
@@ 4.2 Performance on Synthetic Data
Quantitative Results. The results in Tab. 1 demonstrate that OSDFace outperforms other methods in Deg., indicating better alignment with the ID features of HQ images. It also achieves higher scores on LPIPS and DISTS, highlighting superior perceptual coherence and detail preservation. Performance in FID(HQ) shows that OSDFace effectively recovers the distribution of the original HQ dataset.
@@ 4.3 Performance on Real-world Data
Qualitative Results. Figure 7 visualizes representative images in three real-world datasets. Existing face restoration methods often emphasize facial features while overlooking the precise reconstruction of facial accessories like glasses, hair, hats, headwear, and background details. Some methods, especially DiffBIR [32], are over-smoothing and lack fine facial textures. OSDFace naturally and clearly restores multiple faces within a single image, showing its effectiveness for unaligned faces. OSDFace ensures high-fidelity restoration with balanced and consistent visual quality throughout the image. We also test its zero-shot ability on cartoon images in Fig. 6, using Anime Face Dataset [4].
@@ 5 Conclusion
Acknowledgments
We thank the reviewers for their valuable feedback. This work was supported by the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and the Fundamental Research Funds for the Central Universities. Additionally, this work was supported in part by NSFC grant 2024YFC3017100, YDZX20253100004004, 62302299, Shanghai Science and Technology Development Funds under Grant 23YF1420500, and Huawei Explore X funds. Special thanks to Jianze Li for his technical support.